I need help getting started on this problem:

What amount of carbon monoxide occupies a volume of 140 dm3 under normal conditions? (STP)

Since it says ''under normal conditions'', I'm guessing it's a hint that I should use Avogrados Law, but I'm not sure where to begin. I have been given the molecular weight of CO, so should I convert it into moles somehow?

Hoping someone can give me a hint on this, or even a full explanation :-)
1
Expert's answer
2017-03-22T09:49:20-0400
Solution:
Standard conditions for temperature and pressure (STP) are standard sets of conditions for experimental measurements to be established to allow comparisons to be made between different sets of data.
STP is defined as a temperature of 273.15 K (0 °C, 32 °F) and an absolute pressure of exactly 105 Pa (100 kPa, 1 bar).
According to Mendeleeev-Clapeyron’s equation:
pV = nRT, where R is universal gas constant (R = 8.31 J/mol∙K); n is an amount of gas CO.
Thus:
n (CO) = pV/RT
Using given data, we can calculate the n:
n(CO) = (〖10〗^5∙140)/(8.31∙273.15) = 6167.7 (mol).
If we use normal conditions, so the amount of carbon monoxide equals:
n (CO) = (V(CO))/Vm, where Vm is molar volume (Vm = 22.4 L/mol).
That’s why:
n (CO) = 140/22.4 = 6.25 (mol)
Answer: n(CO) = 6.25 moles.